### v1

# Investigating LLM Capabilities on Long Context Comprehension for Medical Question Answering

Abstract
This study is the first to investigate LLM comprehension capabilities over long-context (LC) medical QA of clinical relevance. Our comprehensive assessment spans a range of content-inclusion settings based on their relevance, LLM models of varying capabilities and datasets across task formulations, revealing insights on model size effects, limitations, underlying memorization issues and the benefits of reasoning models. Importantly, we examine the effect of RAG on medical LC comprehension, uncover best settings in single versus multi-document reasoning datasets and showcase RAG strategies for improvements over LC. We shed light into some of the evaluation aspects using a multi-faceted approach. Our qualitative and error analyses address open questions on when RAG is beneficial over LC, revealing common failure cases.
Investigating LLM Capabilities on Long Context Comprehension for Medical Question Answering
Feras AlMannaa1, Talia Tseriotou2, Jenny Chim2, Maria Liakata2,3 1Istanbul Aydın University 2Queen Mary University of London 3The Alan Turing Institute {ferasmhdaymanalmanna}@stu.aydin.edu.tr,{t.tseriotou,m.liakata}@qmul.ac.uk
1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Dubey et al., 2024; Yang et al., 2024) perform impressively on medical tasks (Borgeaud et al., 2021; Nori et al., 2023), achieving superhuman scores on USMLE-style exam question-answering (QA) (Chen et al., 2023; Tang et al., 2023; Pal and Sankarasubbu, 2024; Singhal et al., 2025). Yet multiple-choice QA (MCQA) results are not indicative of performance in more complex tasks Arias-Duart et al. (2025), including open-ended medical QA (Sandeep Nachane et al., 2024) and long-form generation. Furthermore, board exam questions rely primarily on textbook knowledge, likely exposed during pre-training (Chen et al., 2025).
Beyond MCQA, researchers have constructed datasets based on authentic de-identified electronic health records (EHRs). While these benchmarks still face challenges regarding data diversity, ecological, external, and construct validity Wornow et al. (2023), they provide crucial signals on how well models can address complexities in real data, including domain-specific vocabulary, heterogeneous document types, multi-document reasoning, data noise, linguistic diversity, long contexts, and long-range dependencies.
Figure 1: Overview of our Approach
To date most EHR-based benchmarks involve tasks that only require single-document contexts (Pampari et al., 2018; Yue et al., 2020). In practice, there is no guarantee that the necessary information (e.g. to perform a diagnosis) can be found in a single document known a priori; what is needed is comprehension and reasoning over largely noisy and at best complimentary records. EHR-Note-QA (Kweon et al., 2024) is the first to include questions requiring information across records including both multiple-choice (MC) and open-ended questions. There is limited work on the effect of long-range dependencies within biomedical machine reading comprehension (MRC) Vatsal and Singh (2024) and this tends to focus on shorter texts ($\sim$1.5-4K tokens) and biomedical articles and span-based QA rather than complex EHRs. Work that addresses MRC in other domains has utilised methods such as vector compression (Chevalier et al., 2023), memory-based models (Xiao et al., 2024) and context window chunking or sliding (Xiao et al., 2023; Han et al., 2023) to process long context.
Retrieval Augmented Generation (RAG) (Lewis et al., 2021) forms a cost efficient alternative that mitigates risks around information loss or utilizing outdated or incomplete knowledge by selecting relevant context on demand rather than trying to process the entire context. What is preferable is subject to debate, with each showing benefits on different tasks and settings (Li et al., 2024): RAG has been reported to enhance LLM performance compared to long-context prompt compression for QA (Zhang et al., 2024) outperforming SOTA LLMs for both 32K and 128K benchmarks (Xu et al., 2024; Bai et al., 2024). By contrast, others showed that RAG performance peaks at 32K context length for large LLMs (Leng et al., 2024).
Here we investigate how LLMs handle diverse input patient record combinations, using EHR-based human-verified datasets in a variety of long context and RAG settings. Fig. 1 shows a diagram of our study setup. We make the following contributions:
  • We are the first to assess how LLMs of different underlying capabilities and sizes perform on long-context medical QA given longitudinal EHR patient notes. We cover three task formulations: 1) MC , 2) Extractive and 3) Open-ended generative QA. (§3.13.2)
  • We perform comprehensive experiments evaluating the effect of different levels of granularity in the input context while also testing for EHR memorisation (§4.2).
  • We examine the effect of RAG across task formulations and context sizes, showing that given the right setting RAG has superior performance than LC across tasks (§5).
  • We examine the distinct challenges of open-ended QA when reasoning over a single document versus multiple documents (§5).
  • Our qualitative analyses shed light on some of the challenges that LLMs currently face in reasoning over long EHRs (§5.1).
2 Related Work
2.1 Leveraging Long Context
While the increasing long-context window capabilities of some LLMs, i.e. GPT-4o (Hurst et al., 2024) (128K), Claude 3 (Anthropic, 2023) (200K), Gemini 1.5 (Team et al., 2024) (1 million), boost their long-context performance they still struggle in general purpose open-form QA (beyond short passage retrieval) Zhang et al. (2024); Chen et al. (2025). Despite positional embeddings techniques like RoPE (Su et al., 2024), YaRN (Peng et al., 2023) and Position Interpolation (Chen et al., 2023) allowing context extrapolation, QA task performance is low even for 32K context (Chen et al., 2025). Furthermore, the performance gap between open-source and longer-context close-source models (Li et al., 2024) poses questions around the potential avenues for patient-centric long-dependency data where privacy matters.
Benchmarking and assessment of long-context performance of LLMs (Dong et al., 2023; Bai et al., 2024; Liu et al., 2023; Zhang et al., 2024; Hsieh et al., 2024), shows that apart from performance drops with increasing token length LLMs are particularly challenged by long-range dependency QA tasks (Li et al., 2024) that potentially require reasoning. Fan et al. (2024) show notable LLM performance deterioration in medical QA tasks with increased context length up to 200K tokens, with open-source LLMs struggling to produce output given larger contexts. Although long and longitudinal context is prominent in real-world patient and medical data, there is little work investigating LLM capabilities in such settings.
Figure 2: RAG Pipeline
2.2 Retrieval-augmented Generation (RAG)
RAG has been leveraged in a prompt-based plug-in manner to address black-box long-context challenges (Yu et al., 2023). Current retrieval strategies are primarily: chunk-based (Izacard et al., 2021), index-based (Liu, 2022) and summarization-based (Sarthi et al., 2024). LLM-based large-scale retrievers for chunk-based approaches like E5-Mistral-7b (Wang et al., 2022) and BGE M3 (Chen et al., 2024) give notable performance improvement. Furthermore, approaches like Self-RAG (Asai et al., 2023) introduce on-demand retrieval followed by generation and self-reflection outperforming ChatGPT in open-domain QA.
Domain-specific retrievers strongly benefit biomedical retrieval. MedCPT (Jin et al., 2023), a first of its kind contrastive biomedical dense semantic retriever and re-ranker, achieved top performance on 6 biomedical tasks surpassing LLM-based retrievers. BMRETRIEVER (Xu et al., 2024), a biomedical dense retriever trained both contrastively and instruction fine-tuned on synthetic data, shows better performance than MedCPT on most downstream tasks. The Med-RAG (Zhao et al., 2025) method provides better diagnostics over patient EHRs, being enhanced by medical knowledge graph-elicited reasoning. Finally, the Self-BioRAG (Jeong et al., 2024) is a framework that retrieves relevant medical documents using MedCPT and generates through self-reflection on rationales using a domain-specific LLM. Yet Fan et al. (2025) underscore the challenges in retrieval for the medical domain: despite strong medical retrieval capabilities across large corpora, models struggle with specialized medical content such as EHRs. Myers et al. (2025) further demonstrated variability in retrieval across EHR datasets. There is need for work that assesses retrieval performance on patient-oriented QA over long-context.
3 Approach
3.1 Task Selection
To uncover the limitations of LLMs in medical QA given complex patient-oriented clinical scenarios, we focus on content-comprehension settings especially ones where patient context is not supplementary but crucial to complete a task (e.g. obtain diagnoses), thus emphasizing long temporal patient context spanning across time and documents.
Aiming for a comprehensive approach we evaluate datasets covering three task formulations: 1) MCQA, 2) Extractive, and 3) Open-ended QA. We further assess each dataset across four settings with varying note relevance.
3.2 Dataset Selection
Datasets were included in our study based on the following criteria (datasets comparison in Table LABEL:tab:dataset_landscape):
Going beyond Open-domain MCQA: Widely used medical MC benchmarks i.e. MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), PubMedQA (Jin et al., 2019), MMLU (Hendrycks et al., 2021) evaluate option selection instead of grounded synthesis and risk pretraining leakage from exams and textbooks, which can inflate scores, while importantly providing little signal towards noisy clinical settings.
Long-context Documents: To address LLM capabilities in handling long context that reflects real-world settings we seek benchmarks that extend well-beyond 8K. To avoid a simplified needle in the haystack setting we aim for such content to be relevant rather than distracting.
Expert QA Annotations: We selected expert annotated QA pairs grounded on multiple EHR documents to avoid limitations that hinder clinical reliability such as non-expert curation (Fan et al., 2024) or synthetic data (Adams et al., 2024). Patient-centric Longitudinal Content: To study performance on clinically-relevant QA with multi-note reasoning, we prioritize EHR datasets derived from MIMIC-III and MIMIC-IV (Johnson et al., 2016, 2023). These corpora offer standardized de-identified notes across document types (e.g., discharge summaries, radiology reports) and enable consistent retrieval granularity.
3.3 RAG Strategy
RAG strategies can be distinguished into sparse and dense. Sparse retrievers such as BM25 (Robertson et al., 2009) and Splade (Formal et al., 2021) use overlapping terms to match queries with snippets, while dense retrievers encode them into embeddings and match them based on semantic similarity. Recently, using LLM-based embeddings such as E5-Mistral-7B-Instruct (Wang et al., 2023) and Qwen3 Embedding (Zhang et al., 2025) for dense retrieval has been standard practice.
Research on biomedical MCQA (Wang et al., 2024) and biomedical document retrieval (Luo et al., 2022) has shown that hybrid approaches of sparse and dense retrieval outperform single component counterparts. Further, re-ranking of retrieved snippets also exhibits superior performance in biomedical applications (Jin et al., 2023; Wang et al., 2024; Sohn et al., 2025). Additionally, non-hybrid experiments on different EHR subsets of varying granularity (Fan et al., 2025) show that there is no clear performance winner between sparse and dense retrievers. Therefore, we use a hybrid approach of a sparse and denser retriever with a re-ranker.
Our hybrid approach is presented in Figure 2. Clinical notes are segmented into 512-token chunks, which form our document collections $D=\{d_{1},d_{2},\dotsc,d_{n}\}$. The question respectively forms the query $Q$.
For sparse retrieval we employ a lexical retriever as a ranking function of each chunk $d_{j}$ that extends TF-IDF by considering term frequency saturation and document length  [^1] Here terms document and chunk are used interchangeably.:
$S_{\mathrm{sparse}}(Q,d_{j})\;=\;\sum_{i=1}^{n}\mathrm{IDF}(q_{i})\cdot\tfrac{f_{q_{i},d_{j}}\,(k_{1}+1)}{f_{q_{i},d_{j}}+k_{1}\!\left(1-b+b\cdot\frac{|d_{j}|}{\mathrm{avgdl}}\right)}$ (1)
For the dense retriever we use the same Encoder, $E$ to encode the query and each of the chunks and capture semantic similarity of pairs using the cosine similarity:
$S_{\mathrm{dense}}(Q,d_{j})\;=\;\frac{E(Q)\cdot E(d_{j})}{\lVert E(Q)\rVert\,\lVert E(d_{j})\rVert}$ (2)
Top $k$ chunks are retrieved with each retriever and then combined through Reciprocal Rank Fusion (RRF) that takes into account the rank of each document:
$\mathrm{RRF}(d_{j})\;=\;\sum_{s\in\mathcal{S}}\frac{1}{k_{RRF}+\mathrm{rank}_{s}(d_{j})}$ (3)
, where $k_{RRF}$ is a smoothing constant.
Finally, we employ Late Interaction Reranking, a measure of how well every query token is semantically supported by at least one document token, to produce the final top $k$ most contextually relevant set.
$S_{\mathrm{MaxSim}}(Q,d_{j})\;=\;\sum_{i=1}^{|Q|}\;\max_{1\leq r\leq|d_{j}|}\langle\mathbf{q}_{i},\mathbf{d}_{j,r}\rangle$ (4)
Document Ordering: We leverage an order-preserving RAG that sorts retrieved chunks temporally rather than by retrieval score, due to its superior long-context performance (Yu et al., 2024) regardless of the number of chunks.
Note Inclusion: We evaluated two different strategies: direct chunk inclusion, and parent note inclusion (§4.3).
[…]
Table 1: EHR QA Dataset Statistics. Mean/Max Context corresponds to the Include All setting.
3.4 Evaluation Methodology
Evaluating medical QA systems has traditionally relied on automatic metrics and multiple-choice benchmarks Jin et al. (2019, 2021); Pal et al. (2022), which probe memorized knowledge but underrepresent the complexity and factual rigor of clinical reasoning, long-form answers, and EHR-grounded context. To provide clinically meaningful assessment, we adopt a multi-dimensional, reference-based scheme that complements lexical overlap with embedding-based semantic similarity and domain-adapted Natural Language Inference (NLI) for factual consistency and precision. In line with emerging clinical evaluation practices (e.g., MEDIC (Kanithi et al., 2024)) and the rise of LLM-as-a-judge, we employ a calibrated rubric to capture Correctness, Completeness, and Faithfulness at scale, and we prioritize widely available, reproducible metrics while cross-checking signals for robustness. Given that gold standard answers are available, we employ reference-based evaluation:
METEOR Banerjee and Lavie (2005) is used to capture surface-level and lexical similarity.
BERTScore Zhang et al. (2020) computed with Clinical BioBERT embeddings Alsentzer et al. (2019) assesses the semantic similarity.
NLI scores capture the logical relationship between reference and candidate answers. We use a domain-adapted NLI model Deka et al. (2023) with reference answers as premise and candidate answers as hypotheses. We then measure Factual Consistency using the probability of non-contradiction Song et al. (2024) and measure Factual Precision using the entailment probability.
LLM-as-a-judge was employed with a 5-point scale rubric on three aspects, using Qwen-2.5-32B-Instruct as the evaluator model [^2] We selected Qwen for stability and agreement after comparing against Selene-8B (Alexandru et al., 2025) and Prometheus-8x7B v2.0 (Kim et al., 2024) in a small pilot (see Appendix LABEL:sec:judgecomps).. Correctness captures factual consistency with respect to the gold answer, penalizing contradictions while allowing compatible additional information. Completeness captures recall by assessing the predicted answer’s coverage of information present in the reference. Faithfulness captures precision by assessing whether the predicted answer only contains information that is supported by or derivable from the reference.
We report Accuracy on the MC task.
4 Experiments
4.1 Datasets
Context Sources: We leverage MIMIC-III (Johnson et al., 2016) and MIMIC-IV (Johnson et al., 2023) comprising de-identified EHRs covering hospital admissions and ICU stays. They support single- and multi-note settings and long contexts. Our data includes: clinical notes, discharge summaries and radiology reports per patient over time.
QA Datasets: We test on four EHR QA benchmarks spanning different tasks given MIMIC-based underlying context. We only use QA pairs curated or validated by clinical experts to ensure medical validity. We provide a brief dataset overview below, with specifics summarized in Table 1.
CliniQG4QA (Yue et al., 2020): An extractive span benchmark originally created for domain adaptation purposes based on MIMIC-III clinical notes. While the 8,824-QA pair dev set is machine-generated (MG), the 1,287 QA test pairs are expert-annotated using the MG questions for reference. Answers were either generated or verified by three clinical experts. Here we only use the expert-annotated test set.
RadQA (Soni et al., 2022): An extractive QA dataset comprising 3,074 physician-crafted questions and 6,148 answer spans from the Findings and Impressions sections of radiology reports in MIMIC-III. Many questions correspond to multiple sentences and are shown to require different types of reasoning to answer. Annotation is carried out by two human annotators with medical understanding. Here we use the train set and filter out Unanswerable QA pairs resulting in 3,509 pairs.
EHR-DS-QA (Kotschenreuther, 2024): A large synthetic QA corpus of 156,599 pairs generated by two Llama-based (Touvron et al., 2023) LLMs and guided by predefined prompt templates on discharge summaries and notes from MIMIC-IV. A subset of 506 pairs were physician-verified of which we retain 478, marked as correct.
EHRNoteQA (Kweon et al., 2024): Built on MIMIC-IV discharge summaries, it contains 962 QA pairs authored by GPT-4 and iteratively refined by three clinicians. Questions span a diverse set of topics, each corresponding to multiple discharge summaries ($\sim$2.3 per patient). The dataset supports both open-ended and MC formats.
[…]
Table 2: Results for Full Context in each setting across datasets. Bold is best and underlined the second best model in each Setting and Metric. Red highlights the global best across all Settings per Metric. LLM corresponds to LLM Correctness, NLI to NLI Entailment and F1 to BioBERT F1
4.2 Context Formulation
Since most of the datasets are based on a single note/report (CliniQG4QA, RadQA, EHR-DS-QA) or up to three discharge summaries (EHRNoteQA), we identify a large amount of longitudinal patient supplementary content that remains unused. For a more realistic clinical setting, we leverage our context sources, MIMIC-III and MIMIC-IV, to augment few-note context with each patient’s longitudinal notes. We then investigate how LLMs leverage long-form context to answer patient-specific questions. We believe that this setting, (Include All), presents a more realistic scenario where a model would have no prior knowledge of the relevant document and would therefore need to process the entire patient note history [^3] MIMIC preprocessing is described in Appendix C.2..
We explore four data inclusion scenarios depicted in Fig. 1, allowing comprehensive analysis of model sensitivity to context size and noise, ability to handle supplementary information and memorization:
  1. 1.
    Exclude All: Exclusion of all clinical notes - assessing model note memorization.
  2. 2.
    Exclude Relevant: Exclusion of relevant clinical notes - assessing usefulness of supplementary material.
  3. 3.
    Include All: See above.
  4. 4.
    Include Related: Inclusion of only dataset relevant clinical note types: discharge summary(ies), note or radiology report (as marked in each underlying dataset) - assessing model’s ability to leverage relevant context information for patient-centric QA.
Context Segmentation: For the Include All setting we report context-length-based performance across four bins: i) Short context: 0–8K, ii) Medium context: 8–16K, iii) Large context: 16–32K, iv) Extended context: 32–128K tokens.
4.3 Models and Experimental Setup
LLMs: To explore the effect of LLM model size, domain specialization and reasoning capabilities in a controlled way we study four open Qwen2.5 models (Team and others, 2024), namely:
  • Qwen2.5-7B-Instruct
  • HuatuoGPT-o1-Qwen-7B (Chen et al., 2024): medical LLM based on Qwen2.5-7B-Instruct
  • Qwen2.5-32B-Instruct-128K
  • QwQ:32B: reasoning model
Retrieval Settings: We implemented a hybrid retrieval pipeline designed to properly handle the semantic, lexical, and query complexity of medical QA. Our pipeline involves a combination of three different retrievers (more details are in §3.3):
  • Dense Retrieval based on the leading open embedding model Qwen3-Embedding-8B (Zhang et al., 2025).
  • Sparse Retrieval performed using BM25 selected after ablating two commonly used sparse retrievers (Table 8 in Appendix).
  • Late Interaction Reranking with Reason-ModernColBERT, a ColBERT (Khattab and Zaharia, 2020) model finetuned on the ReasonIR dataset (Shao et al., 2025) leading to very high performance on reasoning-intensive retrieval benchmark (Su et al., 2024). Due to its performance intensive nature, we use it as a final reranker in our pipeline.
Experimental Setup: We defer inference specifics to Appendix A and prompt formulation details in Appendix C.3.
5 Results and Discussion
Full context performance for the different settings for all datasets and models is shown in Table 2. For RAG, Table 3 shows different evaluated RAG settings across 3 selected models on the 2 MIMIC-IV based datasets. We excluded QwQ:32B due to time and budget constraints. Similarly, in Table  4 we only evaluate 7B sized models using a single RAG strategy, due to the task nature and limited context size of the MIMIC-III based datasets.
Model Size, Context Size and Performance
Consistent with general findings, larger models consistently perform better across all tasks (Table 2). Qwen2.5:32b and QwQ:32b generally outperformed the 7b parameter models across most metrics and tasks. The performance gap was particularly pronounced in tasks requiring reasoning or synthesizing of information from multiple sources, namely EHRNoteQA and RadQA. However, smaller models performed better in settings excluding relevant information or completely removing the context. This suggests that they are more prone to memorization.
The ‘Include Related’ setting outperforms ‘Include All’ on most datasets and metrics except for the MC task where no benefits from filtering are seen for 3 out of the 4 models. Fig. 3 shows results for Qwen2.5:32B-Instruct over different context sizes. We observed similar trends between other models and context size (see Appendix, Fig.  LABEL:fig:model_performance_context). Worth noting that for context range (8-16K) EHR-DS-QA exhibits a dip in performance across all metrics and models, attributable to noise (see §5.1).
Figure 3: Metric performance over context size on ‘Include All’ for Qwen2.5:32B-Instruct.
Reasoning and Fine-tuning
The impact of reasoning-focused models and medical fine-tuning is mixed. The QwQ:32b general reasoning model showed improved results on Table 2 over other models overall on our two reasoning datasets (RadQA, EHRNoteQA). This suggests that specialized reasoning models can offer an advantage in tasks that demand complex inference across multiple sources of information, but this benefit may be more apparent in larger model sizes where reasoning capabilities can be effectively leveraged without compromising other essential skills.
By contrast, the medically fine-tuned reasoning model, HuatuoGPT-o1-Qwen2.5:7b, did not show any improvement over its standard instruction-tuned base model Qwen2.5:7b-instruct, in line with studies revealing that biomedical LLMs often lead to reduced performance (Dorfner et al., 2024; Dada et al., 2024).
[…]
Table 3: Open-ended and MC QA Results for Context of 8K+ tokens including LC and RAG Methods. Bold is best and underlined the second best performance per Model and Metric. Red highlights the global best across all models per Metric.
RAG Performance
Table 3 provides insights into the effectiveness of different RAG strategies per task and dataset. RAG demonstrated clear performance improvement for the single-note generative task of EHR-DS-QA. In this scenario, the chunk-in-context strategy consistently provided the best results. For the more complex, multi-note reasoning task of EHRNoteQA, the hierarchical RAG strategy was the best with few exceptions. The results also show similar patterns on MC indicating clear performance improvement for all models on the RAG HIR 7 setting. These findings are consistent across different context sizes (Fig.  4).
[…]
Table 4: Extractive QA Results for Context of 4K+ tokens. Bold is best performing per Model and Metric and red highlights the global best per Metric.
Worth noting that using RAG, the medically fine-tuned reasoning model HuatuoGPT-o1-Qwen2.5:7b was able to score the best results on multiple metrics across the 2 datasets. This shows that RAG could benefit smaller models that struggle in FC settings.
While the extractive datasets have shorter contexts overall, Table 4 shows that the RAG setting yields the best performance.
Figure 4: RAG performance across settings with Qwen2.5:32B-Instruct (min-max normalized).
Metrics Insights
Fig. 5 provides insights into the relation between models, metrics, and datasets while focusing on the open-ended generative tasks for the ‘Include All’ and ‘Include Related’ settings combined. Instruct models lead on single-note non-reasoning tasks (EHR-DS-QA). Larger models lead on EHRNoteQA: instruct models are better based on semantic and NLI metrics, while reasoning ones score better on LLM as a Judge metrics  [^4] We include all-context and long-context correlations between metrics in Appendix LABEL:sec:appendix.metric_correlation.
FC vs RAG in QA Tasks
We did a qualitative analysis on the performance of FC and RAG on the generative task across both datasets detailed in Appendix LABEL:sec:appendix.app-rag-vs-fc with examples.
We provide a summary of the main points below:
  • FC performs better for questions requiring comprehensive understanding of the entire document or when answers are located in structured sections, such as summaries, lists, or temporal sequences.
  • RAG outperforms FC for questions that require retrieving specific details or synthesizing information scattered across multiple parts of the document.
  • RAG tends to perform better overall in datasets with complex or lengthy documents, likely due to its ability to focus on relevant information and reduce noise.
  • There is mixed evidence suggesting FC might perform better for tasks involving inferential reasoning or identifying the absence of information.
Figure 5: Model performance across metrics for open-ended QA, normalized with min-max per metric.
5.1 Error Analysis
We analyze failure instances that complement the quantitative trends and provide diagnostic guidance for practical deployment. We summarize three recurrent themes and refer to Appendix LABEL:sec:appendix.case_studies for full cases and diagnostic tables.
Data noise. We observe inconsistencies and ambiguous gold answers in some pairs across all datasets (more profoundly in EHR-DS-QA) that correlate with strange dips in performance at mid-range context range windows. Fan et al. (2024) also observed similar performance dips.
Numerical Values. Questions requiring numeric values (e.g., vitals, dosages, dates) are sensitive to formatting, rounding, or range expressions.
Metric Insights. We did a qualitative analysis to understand the reasons behind metrics’ disagreements across six metrics on the two generative datasets (see details in Appendix LABEL:app:appendix.metric_insights).
We found 233 disagreements in EHRNoteQA and 47 in EHR-DS-QA. These are mainly due to:
  1. 1.
    Correct/Complete but Unfaithful (addition of unsupported details). Most common in EHRNoteQA (173/233; 74%) and present in EHR-DS-QA (17/47; 36%).
  2. 2.
    Correct answer but low surface overlap (lower Meteor/Bert) appears in both datasets (EHRNoteQA: 20%; EHR-DS-QA: 26%).
  3. 3.
    Correct answer but low NLI entailment occurs due to negation/clinical phrasing (EHRNoteQA: 17%; EHR-DS-QA: 21%).
  4. 4.
    EHR-DS-QA shows several short, risk/negation items where NLI flags contradiction despite clinically aligned answers.
This points to the need for thinking about more reasoning appropriate metrics.
6 Conclusion
We studied long-context, patient-centric clinical QA across EHR-grounded datasets, contrasting Full-Context prompting with Retrieval-augmented generation (RAG). Our analysis shows that carefully scoped context (e.g., Include DS) often outperforms feeding all notes, and that hybrid RAG pipelines with reranking are competitive, especially for specific factual queries and multi-note synthesis. Larger models generally help, but performance remains sensitive to context length, evidence placement, and task formulation. Future directions include: (i) scaling beyond single-note or single-source assumptions toward richer multi-note, multi-visit reasoning; (ii) advancing temporal and causal reasoning over longitudinal records, with explicit timeline grounding; (iii) strengthening evaluation via clinically faithful, judge-robust rubrics and cross-metric reliability analyses.
Limitations
Our work focuses on MIMIC-derived datasets, which represents English-only data in a single U.S. medical center. As such, performance may not generalize to other languages, populations, and healthcare systems. Furthermore, the de-identified nature of the data limits us from examining the potential cultural and other biases of LLMs over long contexts. Importantly, although our work presents an early step towards benchmarking long-context LLM performance in longitudinal patient-centric settings, our findings are intended solely for research purposes. Results reflect model performance on question answering benchmarks and should not be interpreted as guarantees of clinical safety, equitable performance, or readiness for clinical deployment.
Beyond complexities of modeling clinical notes, real-world records comprise heterogeneous data, including data from other sources and in other modalities. Our work poses limitations in the assessment of long-context by focusing only on the textual modality, something we aim to address in future work.
Despite the in depth study of literature and our initial ablations in selecting a competitive RAG methodology, our work does not exhaustively examine the full potential of RAG under different settings, i.e. different retrievers and chunk sizes, in long-form medical QA. Finally, while we focus on Qwen-based models due to their competitive performance, we acknowledge that an evaluation of a wider range of LLMs could offer a more complete picture of the LLM landscape in long-form medical QA. Our analysis has also pointed to the limitations of current evaluation metrics, both in terms of assessing reasoning as well as lack of transparency in LLM as a judge metrics and how faithfulness can be truly assessed currently when additional information, not present in the benchmark, is added.
Ethics Statement
This work uses MIMIC-III, MIMIC-IV, and four derivative datasets (EHRNoteQA, EHR-DS-QA, RadQA, CliniQG4QA) accessed under the PhysioNet Credentialed Health Data License 1.5.0. All datasets contain de-identified patient records in accordance with HIPAA Safe Harbor standards. We exclusively used open-weight LLMs to ensure no patient data were transmitted to third-party proprietary systems.
Appendix A Inference Setup
  • Inference Engine: Nvidia TensorRT.
  • Quantization: FP8.
  • Hardware: 7B parameter models: Deployed on single Nvidia H100 GPUs with 40GB VRAM. 32B parameter models with extended context: Deployed on dual Nvidia H100 GPUs, each equipped with 80GB VRAM. A total of 260 GPU hours was spent.
  • Libraries: Qdrant, DSPy, Litellm, pandas, Evaluate, Fastembed, Pylate
Appendix B Hyperparameters
Hyperparameter Value
LLM temperature inst: 0, reas: 11
freq_penalty 0
pres_penalty 0
think tokens QwQ: $20k$
HuatuoGPT: $8k$
RAG top_k $2\times$ num chunks
$k_{\text{RRF}}$ 6060
$k_{1}$ TF saturation
$b$ length norm.
avgdl avg. doc length
Table 5: Hyperparameter Configuration
Appendix C Datasets
C.1 Context Data Distributions
Token-length distributions per dataset are shown in Figures 6 and 7.
C.2 Context Data Preparation
We performed minimal cleaning and standardization to support retrieval and prompting, including merging notes by patient and stay, normalizing timestamps, adding note-type metadata, computing token counts for context segmentation, and filtering to human-verified subsets where applicable.
We describe the cleaning and normalization steps applied prior to indexing and prompting:
  • Cleaning and de-duplication of clinical notes; removal of template boilerplate where applicable.
  • Merging notes by patient and hospital stay; preserving note-type and section headers for downstream retrieval.
  • Datetime standardization and chronological ordering; normalization where timezones or partial timestamps occur.
  • Metadata augmentation (e.g., note type, encounter identifiers) to support Include DS vs Include All settings.
  • Tokenization and token-count computation at note- and patient-level for context segmentation.
Figure 6: MIMIC-III Context Distributions
Figure 7: MIMIC-IV Context Distributions
These details enable reproducibility of sampling and retrieval segmentation.
Dataset LLM-as-a-judge QA Pairs EHR-DS-QA Selene-8B 113 Prometheus-8x7B-v2.0 98 Qwen2.5-32B-Instruct 71 EHRNoteQA Selene-8B 180 Prometheus-8x7B-v2.0 220 Qwen2.5-32B-Instruct 42
Table 6: Disagreement instances where NLI Med Contradiction is high (>> 0.7) and LLM Correctness is also high (>> 0.7).
Dataset LLM-as-a-judge 1 LLM-as-a-judge 2 QA Pairs EHR-DS-QA Prometheus Qwen2.5-32B-Instruct 520 Prometheus Selene-8B 363 Qwen2.5-32B-Instruct Selene-8B 664 EHRNoteQA Prometheus Qwen2.5-32B-Instruct 690 Prometheus Selene-8B 236 Qwen2.5-32B-Instruct Selene-8B 519
Table 7: Judge-pair disagreements where the absolute difference in LLM Correctness $\geq$ 0.5.
Metric Splade BM25 LLM Correctness 67.30 69.34 LLM Completeness 65.41 67.14 LLM Faithfulness 47.64 46.23 NLI Med Entailment 52.68 54.35 NLI Med Contradiction 16.98 16.94 Bio BERTScore F1 79.75 79.67 METEOR 43.93 43.63
Table 8: Comparison of Different Sparse Retrievers
C.3 Prompt Formulations
We used a simple one shot prompt structure and tailored it explicitly for each task formulation:
  • Extractive: Request the model to answer by extracting the most relevant answer from the context.
  • Multiple-choice: Standard multiple choice prompt.
  • Open-ended: Focusing on open ended question answering favoring short, single sentence answers.
Below is the sample prompt for extractive tasks:
System message: Your input fields are: 1. medical_record (list[str]): List of patient notes (chronological order). 2. question (str): A question about the patients record. Your output fields are: 1. answer (str): Short single-sentence answer to the question. All interactions will be structured in the following way, with the appropriate values filled in. [[ ## medical_record ## ]] {medical_record} [[ ## question ## ]] {question} [[ ## answer ## ]] {answer} [[ ## completed ## ]] In adhering to this structure, your objective is: Given a patients medical record and a question, answer the question correctly and shortly.
C.4 Dataset Comparisons