### v2

# Investigating LLM Capabilities on Long Context Comprehension for Medical Question Answering

Abstract
This study is the first to investigate LLM comprehension capabilities over long-context (LC), clinically relevant medical Question Answering (QA) beyond MCQA. Our comprehensive approach considers a range of settings based on content inclusion of varying size and relevance, LLM models of different capabilities and a variety of datasets across task formulations. We reveal insights on model size effects and their limitations, underlying memorization issues and the benefits of reasoning models, while demonstrating the value and challenges of leveraging the full long patient’s context. Importantly, we examine the effect of Retrieval Augmented Generation (RAG) on medical LC comprehension, showcasing best settings in single versus multi-document QA datasets. We shed light into some of the evaluation aspects using a multi-faceted approach uncovering common metric challenges. Our quantitative analysis reveals challenging cases where RAG excels while still showing limitations in cases requiring temporal reasoning.
Investigating LLM Capabilities on Long Context Comprehension for Medical Question Answering
Feras AlMannaa1∗, Talia Tseriotou2∗, Jenny Chim2, Maria Liakata2,3 1Istanbul Aydın University 2Queen Mary University of London 3The Alan Turing Institute {ferasmhdaymanalmanna}@stu.aydin.edu.tr,{t.tseriotou,m.liakata}@qmul.ac.uk
* Indicates equal contribution.
1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Grattafiori et al., 2024; Yang et al., 2024) perform impressively on medical tasks (Borgeaud et al., 2022; Nori et al., 2023), achieving superhuman scores on United States Medical Licensing Examination (USMLE)-style exam question-answering (QA) (Chen et al., 2023b; Tang et al., 2024; Pal and Sankarasubbu, 2024; Singhal et al., 2025). Yet evaluation is mostly based on multiple-choice QA (MCQA) (Liévin et al., 2024; Xiong et al., 2024; Chen et al., 2024b; Singhal et al., 2025) which doesn’t reflect performance in complex tasks Arias-Duart et al. (2025), such as open-ended medical QA (Sandeep Nachane et al., 2024). Furthermore, medical board exams designed to assess professional knowledge and decision-making rely primarily on textbook knowledge which is likely available during pre-training (Chen et al., 2025a). Assessing LLM capabilities on context from expert-curated electronic health records (EHRs) could provide crucial signals on how well models address complexities in real data, including domain-specific vocabulary, heterogeneous document types, multi-document reasoning, data noise, linguistic diversity, long contexts, and long-range dependencies Wornow et al. (2023). However, research on open-form QA pertaining to EHRs focuses on creating datasets (Yang et al., 2022) or EHR-specialized models (Fleming et al., 2024), rather than investigating LLM performance in real-world QA.
Figure 1: Overview of our Approach
Moreover, most EHR-based benchmarks consider single-document contexts (Pampari et al., 2018; Yue et al., 2021), while in practice there is no guarantee that the necessary information will be included within a specific patient document. Reasoning over multiple long, temporally-dependent documents further poses questions around long-context (LC) limitations and evidence positioning sensitivity (Liu et al., 2024; Xiao et al., 2023), also observed in medically focused studies (Adams et al., 2025). Work on the effect of long-range dependencies within medical machine reading comprehension (MRC) is limited Vatsal and Singh (2024), focusing on shorter texts ($\sim$1.5-4K tokens) and span-based QA rather than complex EHRs. Work on LC medical QA involves either synthetic data, is MCQA-based and still relatively short (up to 6k) (Adams et al., 2025)), or small in sample size and not human-validated (Fan et al., 2025a). We address this gap by investigating LLM capabilities to answer questions given longitudinal patient-centric EHRs, where QA pairs are human-validated.
While recent LLM releases push the frontiers of context size, performance does not scale with increased context (Levy et al., 2024; Modarressi et al., 2025). Retrieval Augmented Generation (RAG) (Lewis et al., 2020) often forms a cost efficient alternative by selecting relevant context on demand. Whether LC or RAG is preferable is subject to debate, with each showing benefits for different tasks and settings (Li et al., 2024b). RAG has been reported to enhance LLM performance compared to LC prompt compression for QA (Zhang et al., 2024a), outperforming SOTA LLMs on both 32K and 128K benchmarks (Xu et al., 2024a; Bai et al., 2025). By contrast, others showed that RAG performance peaks at 32K context length for large LLMs (Leng et al., 2024). While work on medical QA reports promising gains when retrieval is carefully designed (Xiong et al., 2024), LC vs RAG superiority in long-form medical QA remains an open question. We make the following contributions:
  • We are the first to assess how LLMs of different underlying capabilities and sizes perform on long-context medical QA given longitudinal EHR patient notes, covering three task formulations: 1) MCQA , 2) Extractive and 3) Open-ended generative QA (§4.14.3) and a variety of LC and RAG settings, as depicted in Fig. 1.
  • We perform comprehensive experiments evaluating different note inclusion strategies in the input context, including note memorization (§4.2).
  • We carefully design a hybrid RAG pipeline examining its performance over Full-context (FC) across task formulations and context sizes, showing its superiority across tasks (§5.15.2).
  • Our experiments and multi-faceted evaluation methodology allow for a direct comparison between long-form reasoning when the answer is part of a single versus multiple documents (§5.1).
  • Our quantitative and qualitative analyses uncover some of the LLM and metrics challenges in reasoning over long EHRs, while demonstrating such cases where RAG excels (§5.25.3).
2 Related Work
2.1 Leveraging Long Context
While increasing long-context window capabilities of LLMs, e.g. GPT-4o (Hurst et al., 2024) (128K), Claude 3 (Anthropic, 2023) (200K), Gemini 1.5 (Team et al., 2024a) (1 million), boost their long-context performance they still struggle in general purpose open-form QA (beyond short passage retrieval) Zhang et al. (2024b); Chen et al. (2025b). Despite positional embedding techniques like RoPE (Su et al., 2024), YaRN (Peng et al., 2023) and Position Interpolation (Chen et al., 2023a) allowing context extrapolation, QA task performance is low even for 32K contexts (Chen et al., 2025b). Furthermore, the performance gap between open-source and longer-context close-source models (Li et al., 2024a) poses questions around the potential avenues for patient-centric long-dependency data where privacy matters.
Assessment of long-context LLM performance (Dong et al., 2024; Bai et al., 2024; Liu et al., 2024; Zhang et al., 2024b; Hsieh et al., 2024), shows that apart from performance drops with increasing token length, LLMs are particularly challenged by long-range dependencies (Li et al., 2024a) that potentially require reasoning. Fan et al. (2025a) show notable LLM performance drop in medical QA tasks with contexts up to 200K tokens, with open-source LLMs struggling to produce output given larger contexts. Although long and longitudinal context is prominent in real-world patient and medical data, there is little work investigating LLM capabilities in such settings.
Figure 2: RAG Pipeline
2.2 Retrieval-augmented Generation (RAG)
RAG has been leveraged in a prompt-based plug-in manner to address black-box long-context challenges (Yu et al., 2023). Current retrieval strategies are primarily: chunk-based (Izacard et al., 2021), index-based (Liu, 2022) and summarization-based (Sarthi et al., 2024). LLM-based large-scale retrievers for chunk-based approaches like BGE M3 (Chen et al., 2024a) give notable performance improvements. Furthermore, approaches like Self-RAG (Asai et al., 2024) outperform ChatGPT in open-domain QA through on-demand retrieval followed by generation and self-reflection.
Biomedical retrievers showed benefits in domain retrieval, with the MedCPT dense semantic retriever and re-ranker (Jin et al., 2023), achieving top performance on 6 biomedical tasks surpassing LLM-based counterparts. The BMRETRIEVER (Xu et al., 2024b) further surpasses MedCPT’s performance on most downstream tasks. More recently, methods like Med-RAG (Zhao et al., 2025) showed better diagnostics over EHRs, through enhancement by knowledge graph-elicited reasoning, while Self-BioRAG (Jeong et al., 2024) employs MedCPT and generates through self-reflection using a domain-specific LLM. Yet Fan et al. (2025b) underscore the medical retrieval challenges despite strong in-domain retrieval capabilities across large corpora, with models struggling with specialized medical content such as EHRs. Myers et al. (2025) further demonstrate variability in retrieval across EHR datasets. Therefore a thorough examination of RAG vs LC over realistic long-context patient-oriented QA is needed.
3 Approach
Since strong medical QA performance offers the potential of LLM integration in clinical workflows we aim to uncover LLM limitations given complex patient-oriented scenarios. We focus on content-comprehension settings using longitudinal EHRs, thus emphasizing relevant patient-centric long-context across time and documents.
Our comprehensive approach spans across expert annotated datasets covering three task formulations: 1) MCQA, 2) Extractive, and 3) Open-ended QA. We assess each dataset across four settings with varying note relevance (see §4.2).
3.1 Dataset Selection Criteria
The following criteria were used for benchmark selection in our study (full comparison in Table LABEL:tab:dataset_landscape):
Beyond MCQA: Medical MCQA benchmarks i.e. PubMedQA (Jin et al., 2019), MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022) evaluate option selection instead of grounded synthesis and risk pretraining leakage from textbooks, which can inflate scores, providing little signal towards noisy clinical settings. Our evaluation goes beyond MCQA to both extractive and open-ended QA.
Long-context Documents: To address LLM capabilities in handling real-world LC we seek benchmarks extending well-beyond 8K, with inclusion of patient-specific content.
Expert Annotated QA pairs grounded on multi-document EHRs were our choice, to avoid clinical reliability issues like non-expert curation (Fan et al., 2025a) or synthetic data (Adams et al., 2025).
Patient-centric Longitudinal Content: Prioritizing clinically-relevant QA with reasoning, we focus on MIMIC-based EHR datasets (Johnson et al., 2016, 2023), offering timestamped notes across document types with consistent format.
3.2 RAG Approach
Retrievers are distinguished into sparse and dense. Sparse retrievers such as BM25 (Robertson and Zaragoza, 2009) and Splade (Formal et al., 2021) use overlapping terms to match queries with snippets, while dense retrievers encode them into embeddings and match them based on semantic similarity. Recently, using LLM-based embeddings for dense retrieval has become standard practice.
Research on biomedical MCQA (Wang et al., 2024) and biomedical document retrieval (Luo et al., 2022) has shown that hybrid approaches combining sparse and dense retrieval outperform single component counterparts. Additionally, non-hybrid experiments on EHR subsets of varying granularity (Fan et al., 2025b) show no clear winner between sparse and dense retrievers. Finally, re-ranking of retrieved snippets also exhibits superior performance in biomedical applications (Jin et al., 2023; Wang et al., 2024; Sohn et al., 2025). Based on the above, we leverage a hybrid approach followed by re-ranking, presented in Figure 2. Clinical notes are segmented into 512-token chunks, forming the document collections $D=\{d_{1},d_{2},\dotsc,d_{n}\}$. The question respectively forms the query $Q$.
Sparse retrieval: we employ a lexical retriever as a ranking function of each chunk $d_{j}$  [^1] Here terms document and chunk are used interchangeably.:
$S_{\mathrm{sparse}}(Q,d_{j})\;=\;\sum_{i=1}^{n}\mathrm{IDF}(q_{i})\cdot\tfrac{f_{q_{i},d_{j}}\,(k_{1}+1)}{f_{q_{i},d_{j}}+k_{1}\!\left(1-b+b\cdot\frac{|d_{j}|}{\mathrm{avgdl}}\right)}$ (1)
Dense retrieval: we use the same encoder, $E$, for the query and the chunks and capture semantic similarity of such pairs using the cosine similarity:
$S_{\mathrm{dense}}(Q,d_{j})\;=\;\frac{E(Q)\cdot E(d_{j})}{\lVert E(Q)\rVert\,\lVert E(d_{j})\rVert}$ (2)
Reciprocal Rank Fusion (RRF): combines the top-$k$ chunks from each retriever accounting for each document’s rank, using a smoothing constant $k_{RRF}$:
$\mathrm{RRF}(d_{j})\;=\;\sum_{s\in\mathcal{S}}\frac{1}{k_{RRF}+\mathrm{rank}_{s}(d_{j})}$ (3)
Late Interaction Reranking: measures how well every $Q$-token is semantically supported by at least one document token, to produce the final top-$k$ set:
$S_{\mathrm{MaxSim}}(Q,d_{j})\;=\;\sum_{i=1}^{|Q|}\;\max_{1\leq r\leq|d_{j}|}\langle\mathbf{q}_{i},\mathbf{d}_{j,r}\rangle$ (4)
Document Ordering: we sort retrieved chunks temporally rather than by retrieval score due to its superior long-context performance (Yu et al., 2024).
[…]
Table 1: EHR QA Dataset Statistics. Mean/Max Context corresponds to the Include All setting.
3.3 Evaluation Methodology
Evaluating medical QA systems has traditionally relied on automatic metrics and MCQA benchmarks Jin et al. (2019, 2021); Pal et al. (2022), which probe memorized knowledge and underrepresent the complexity of clinical reasoning, and EHR-grounded context. Towards a clinically meaningful assessment, we adopt a multi-dimensional, reference-based scheme that complements lexical overlap with embedding-based semantic similarity and domain-adapted Natural Language Inference (NLI). In line with emerging clinical evaluation practices (e.g., MEDIC (Kanithi et al., 2024)) and the rise of LLM-as-a-judge, we employ a calibrated rubric and we prioritize widely available metrics while cross-checking signals for robustness. Our datasets comprise gold-standard QA pairs that are either expert-verified or expert-generated enabling a clinically centric evaluation:
METEOR Banerjee and Lavie (2005) to capture surface-level and lexical similarity.
BERTScore Zhang et al. (2019), computed with Clinical BioBERT embeddings Alsentzer et al. (2019) to assess semantic similarity.
NLI Scores, to capture the logical relationship between reference and candidate answers. We use a domain-adapted NLI model Deka et al. (2023) and then measure Factual Consistency using the probability of non-contradiction Song et al. (2024) and Factual Precision using the entailment probability.
LLM-as-a-judge, employed with a 5-point scale rubric on three aspects: Correctness capturing factual consistency with respect to the gold answer, penalizing contradictions while allowing compatible additional information, Completeness capturing recall by assessing the predicted answer’s coverage of information present in the reference, Faithfulness capturing precision assessing if the predicted answer only contains information that is supported by the reference.
Accuracy reported on the MCQA task.
4 Experiments
4.1 Datasets
We leverage MIMIC-III (Johnson et al., 2016) and MIMIC-IV (Johnson et al., 2023), comprising de-identified EHRs covering hospital admissions and ICU stays, as the underlying context. They support single- and multi-note settings and include: clinical notes (CN), discharge summaries (DS) and radiology reports (RR) per patient over time.
Our four selected QA datasets only use QA pairs curated or validated by clinical experts to ensure medical validity. We provide a brief dataset overview below, also summarized in Table 1:
CliniQG4QA (Yue et al., 2021): An extractive span benchmark based on MIMIC-III. While most of its pairs are machine-generated (MG), the 1,287 QA test pairs are either generated or verified by three clinical experts using the MG questions for reference. We only use the expert-annotated set.
RadQA (Soni et al., 2022): An extractive dataset comprising 3,074 physician-crafted questions and 6,148 answer spans from the Findings and Impressions sections of RR in MIMIC-III. Many questions correspond to multiple sentences and require different types of reasoning. Annotation is carried out by two human annotators. After filtering out for Unanswerable pairs we obtain 3,509 pairs.
EHR-DS-QA (Kotschenreuther, 2024): A synthetic QA corpus of 156,599 pairs generated by two LLMs and guided by predefined prompt templates on DS and CN from MIMIC-IV. A subset of 506 pairs were physician-verified of which we retain 478, marked as correct.
EHRNoteQA (Kweon et al., 2024): Built on MIMIC-IV, it contains 962 QA pairs authored by GPT-4 and iteratively refined by three clinicians. Questions span a diverse set of topics, each corresponding to multiple DS ($\sim$2.3 per patient), while supporting both open-ended and MCQA formats.
[…]
Table 2: Results for Full Context in each setting across datasets. Bold is best and underlined the second best model in each Setting and Metric. Red highlights the global best across all Settings per Metric. LLM corresponds to LLM Correctness, NLI to NLI Entailment and F1 to BioBERT F1
4.2 Context Formulation
While most of the datasets are based on a single note/report (CliniQG4QA, RadQA, EHR-DS-QA) or up to three DS (EHRNoteQA), there is a large amount of longitudinal patient supplementary content that remains unused. Aiming for a realistic clinical setting where there is no a-priori knowledge of the most relevant note, we leverage the MIMIC resources to augment the context with each patient’s longitudinal notes [^2] MIMIC preprocessing is described in Appendix C.2., resulting in long context from the entire patient’s note history (Include All). We then investigate how LLMs leverage long-form context to answer patient-specific questions.
We explore four data inclusion scenarios depicted in Fig. 1, allowing comprehensive analysis:
  1. 1.
    Exclude All: Exclusion of all notes - assessing model note memorization.
  2. 2.
    Exclude Relevant: Exclusion of relevant notes - assessing usefulness of supplementary material.
  3. 3.
    Include All: Inclusion of all patient notes - assessing LC processing of varying importance.
  4. 4.
    Include Related: Inclusion of only the most relevant note(s) (DS, CN or RR) as marked in each dataset - assessing model’s ability to leverage the most relevant context information.
For the Include All setting we report context-length-based performance across four token bins: Short context: 0–8K, Medium context: 8–16K, Large context: 16–32K, Extended context: 32–128K.
4.3 Models and Experimental Setup
LLMs: To explore the effect of model size, domain specialization and reasoning capabilities we study four open Qwen2.5 (Team et al., 2024b) single-family models, ensuring a controlled comparison:
  • Qwen2.5-7B-Instruct
  • HuatuoGPT-o1-Qwen-7B (Chen et al., 2024b): medical LLM based on Qwen2.5-7B-Instruct
  • Qwen2.5-32B-Instruct-128K
  • QwQ:32B: reasoning model
Evaluation: Through comparisons against Selene-8B (Alexandru et al., 2025) and Prometheus-8x7B v2.0 (Kim et al., 2024), we finally selected Qwen-2.5-32B-Instruct for the LLM-as-a-judge for stability and agreement (see Appendix LABEL:sec:judgecomps).
Retrieval Settings: Based on biomedical QA retrieval research that demonstrates the superiority of combining sparse and dense retrievers (see §3.2), we designed a hybrid retrieval approach to properly handle the semantic, lexical, and query complexity. The selected model components are enlisted below:
  • Dense Retrieval using the open embedding model Qwen3-Embedding-8B (Zhang et al., 2025).
  • Sparse Retrieval using BM25, selected after ablating two sparse retrievers (see Table 9).
  • Late Interaction Reranking with Reason-ModernColBERT, a ColBERT (Khattab and Zaharia, 2020) model finetuned on the ReasonIR dataset (Shao et al., 2025) demonstrating high performance on a reasoning-intensive retrieval benchmark (Su et al., 2025).
We explored two chunk inclusion strategies: i) direct chunk inclusion (RAG), ii) hierarchical parent note inclusion (RAG HIR).
Experimental Setup: LLM inference and prompt formulation specifics are in Appendices A and C.3.
5 Results and Discussion
5.1 Main Results
Performance for the different inclusion settings across datasets and models is shown in Table 2.
Model Size, Tasks and Inclusion Settings:
Consistent with general findings, larger models perform better across all tasks (Table 2). Qwen2.5:32B and QwQ:32B generally outperform the 7B parameter models across most metrics and tasks. The performance gap is particularly pronounced in tasks requiring reasoning or information synthesis from multiple sources, namely EHRNoteQA and RadQA. However, smaller models perform better in settings excluding relevant information or completely removing the context. This suggests that they are more prone to memorization. Additionally, unlike the rest of the tasks, MC shows strong evidence of memorization with a performance of 59.24% in the ‘Exclude All’ setting. While MC demonstrates 90%+ performance, open-ended and extractive tasks remain challenging for LLMs.
The ‘Include Related’ setting outperforms ‘Include All’ on most datasets and metrics except for MC. Despite that, the ‘Exclude Related’ setting outperforms the ‘Exclude All’ by a margin across the board demonstrating the usefulness of the supplementary patient material and suggesting that while the ‘Include All’ setting includes useful additional information, LLMs struggle to process long context and to identify the most relevant information.
Figure 3: Metric performance over context size on ‘Include All’ for Qwen2.5:32B-Instruct.
Reasoning and Fine-tuning:
The impact of reasoning-focused models and medical fine-tuning is mixed. The QwQ:32B general reasoning model showed improved results on Table 2 over other models overall on RadQA, EHRNoteQA - the reasoning datasets. This suggests that reasoning models can offer an advantage in tasks that demand complex inference across multiple sources of information, but this benefit may be more apparent in larger model sizes where reasoning capabilities can be effectively leveraged without compromising other essential skills. By contrast, the medically fine-tuned reasoning model, HuatuoGPT-o1-Qwen2.5:7B, did not show any improvement over its standard instruction-tuned base model Qwen2.5:7B-instruct, in line with studies revealing that biomedical LLMs often lead to reduced performance (Dorfner et al., 2024; Dada et al., 2025).
[…]
Table 3: Open-ended and MCQA Results for Context of 8K+ tokens including LC and RAG Methods. Bold is best and underlined the second best performance per Model and Metric. Red highlights the global best across all models per Metric.
[…]
Table 4: Extractive QA Results for Context of 4K+ tokens. Bold is best performing per Model and Metric and red highlights the global best per Metric.
Context Size:
Fig. 3 shows results for Qwen2.5:32B-Instruct over different context sizes. Generally, performance decreases across metrics with increased context size on open-ended QA particularly when multi-note reasoning is required (EHRNoteQA), showing that even large LLMs are challenged by long context, with similar trends observed across models (see Appendix, Fig.  LABEL:fig:model_performance_context). Worth noting that for context range (8-16K) EHR-DS-QA exhibits a dip in performance across all metrics and models. This is likely attributed to data noise, based on manually observed inconsistencies and ambiguous gold answers in some pairs especially in EHR-DS-QA, correlating with strange dips in performance at mid-range context windows. Fan et al. (2025a) and Ma et al. (2025) observed similar dips.
Figure 4: RAG performance across settings with Qwen2.5:32B-Instruct (min-max normalized).
RAG Performance:
Table 3 shows the two evaluated chunk settings on the open-ended generative datasets [^3] We excluded QwQ:32B due to time and budget constraints. RAG demonstrated clear performance improvement for the single-note generative task of EHR-DS-QA. In this scenario, the direct chunk inclusion strategy consistently provided the best results. For the more complex, multi-note reasoning task of EHRNoteQA, the hierarchical RAG strategy (RAG HIR) was the best with few exceptions. The results also show similar patterns on MC indicating clear performance improvement for all models on the RAG HIR 7 setting. Our findings are consistent across medium and larger (8K+) context sizes (Fig.  4). Worth noting that using RAG, the medically fine-tuned reasoning model HuatuoGPT-o1-Qwen2.5:7B scored the best results on multiple metrics across the open-ended datasets, showing that RAG could benefit even more so the smaller models that struggle in Full-context (FC) settings (‘Include All’).
In Table  4 we only evaluate 7B sized models using a single RAG setting, due to the task nature and limited context size of the extractive datasets. While the extractive datasets have shorter contexts overall RAG still yields the best performance.
Category Total Favored by RAG Favored by FC SEE 18 12 (66.7%) 6 (33.3%) TLR 10 3 (30.0%) 7 (70.0%) CNS 10 6 (60.0%) 4 (40.0%) ANI 3 2 (66.7%) 1 (33.3%) Total 41 23 (56.1%) 18 (43.9%)
Table 5: Breakdown of sampled favored by RAG vs FC across four categories in open-ended QA.
5.2 Quantitative Analysis: RAG vs FC
In our analysis we gathered challenging cases of disagreement between metrics (LLM Correctness, NLI Entailment, and Bio-F1) across EHRNoteQA and EHR-DS-QA. Metrics were first z-score standardized and samples were gathered for cases were some metrics belonged to the top 60th percentile while another metric fell within the bottom 40th percentile. The 41 examples gathered were then manually examined in order to categorize them across QA types and determine which answer between RAG and FC is more reliable. The rubric is provided below:
  • Specific Entity Extraction (SEE): Requiring retrieval of high cardinality facts such as medications and lab values or surgical procedures.
  • Temporal & Longitudinal Reasoning (TLR): Requiring tracking of changes over time, chronological event ordering or visit comparisons.
  • Clinical Nuance & Status (CNS): Questions regarding discharge instructions, patient mental status, or synthesized clinical course summaries.
  • Absence or Negative Information (ANI): Requiring the identification of absent information.
Figure 5: Model performance across metrics for open-ended QA, normalized with min-max per metric.
In Table 5, RAG shows higher performance across the SEE, CNS and ANI categories, while FC leads in the TLR one, demonstrating that while RAG effectively preserves the relevant information for challenging cases, it falls behind when temporal reasoning is required. A further qualitative analysis providing insights on cases where RAG or FC perform better (see Appendix LABEL:sec:appendix.app-rag-vs-fc) showcases findings on par with the quantitative ones, with the exception of ANI where FC has more promising although inconsistent performance.
5.3 Metrics and Insights
Metric Comparisons: Fig. 5 provides insights into the relation between models, metrics, and datasets while focusing on the open-ended generative tasks for the ‘Include All’ and ‘Include Related’ settings combined. Instruct models lead on the single-note non-reasoning task (EHR-DS-QA), while larger models lead on the multi-note reasoning one (EHRNoteQA). Instruct models are better based on semantic and NLI metrics, while reasoning ones score better on LLM-as-a-judge metrics [^4] We include all-context and long-context correlations between metrics in Appendix LABEL:sec:appendix.metric_correlation.
Metric Insights: We performed a qualitative analysis to investigate the reasons behind metrics’ disagreements on our open-ended datasets (see Appendix LABEL:app:appendix.metric_insights). The 233 disagreements in EHRNoteQA and 47 in EHR-DS-QA are mainly due to:
  • High Correctness/Completeness but low Faithfulness in predictions due to the addition of unsupported details - in both datasets but more pronounced in EHR-DS-QA.
  • High Correctness but with low surface overlap between the gold and the prediction (lower METEOR/BERTScore) - in both datasets.
  • High Correctness but with either low NLI entailment or high NLI contradiction occurring due to either existence of negation or shortness in prediction or differences in clinical phrasing - in both datasets.
These point to the need for thinking about more reasoning appropriate metrics [^5] Manually analyzed failure cases are in Appendix LABEL:sec:appendix.case_studies..
6 Conclusion
We studied long-context, patient-centric clinical QA across EHR-grounded datasets, contrasting Full-Context (FC) prompting with Retrieval-Augmented generation (RAG). Our work shows that highly-relevant context often outperforms feeding all notes, likely due to LLMs struggling with long context. Hybrid RAG pipelines with reranking overperform FC, especially for specific factual queries and multi-note synthesis. Larger models generally help especially for reasoning datasets, but performance remains sensitive to context length and task formulation, while medically fine-tuned models fall behind. Future directions include: (i) scaling beyond single-note or single-source assumptions toward richer multi-note, multi-visit reasoning; (ii) advancing temporal and causal reasoning over longitudinal records, with explicit timeline grounding; (iii) strengthening evaluation via clinically faithful, judge-robust rubrics and cross-metric reliability analyses.
Limitations
Our work focuses on MIMIC-derived datasets, which represents English-only data in a single U.S. medical center. As such, performance may not generalize to other languages, populations, and healthcare systems. Furthermore, the de-identified nature of the data limits us from examining the potential cultural and other biases of LLMs over long contexts. Importantly, although our work presents an early step towards benchmarking long-context LLM performance in longitudinal patient-centric settings, our findings are intended solely for research purposes. Results reflect model performance on question answering benchmarks and should not be interpreted as guarantees of clinical safety, equitable performance, or readiness for clinical deployment.
Beyond complexities of modeling clinical notes, real-world records comprise heterogeneous data, including data from other sources and in other modalities. Our work poses limitations in the assessment of long-context by focusing only on the textual modality, something we aim to address in future work.
Despite the in-depth study of literature and our initial ablations in selecting a competitive RAG methodology, our work does not exhaustively examine the full potential of RAG under different settings, i.e. different retrievers and chunk sizes, in long-form medical QA. Finally, while we focus on Qwen-based models due to their competitive performance, we acknowledge that an evaluation of a wider range of LLMs could offer a more complete picture of the LLM landscape in long-form medical QA. Our analysis also highlights limitations in current evaluation metrics, specifically regarding the assessment of reasoning, the lack of transparency in LLM-as-a-judge frameworks, and the difficulty of accurately measuring faithfulness when benchmarks are supplemented with additional information.
Ethics Statement
This work uses MIMIC-III, MIMIC-IV, and four derivative datasets (EHRNoteQA, EHR-DS-QA, RadQA, CliniQG4QA) accessed under the PhysioNet Credentialed Health Data License 1.5.0. All datasets contain de-identified patient records in accordance with HIPAA Safe Harbor standards. We exclusively used open-weight LLMs to ensure no patient data were transmitted to third-party proprietary systems.
Appendix A Inference Setup
  • Inference Engine: Nvidia TensorRT.
  • Quantization: FP8.
  • Hardware: 7B parameter models: Deployed on single Nvidia H100 GPUs with 40GB VRAM. 32B parameter models with extended context: Deployed on dual Nvidia H100 GPUs, each equipped with 80GB VRAM. A total of 260 GPU hours was spent.
  • Libraries: Qdrant, DSPy, Litellm, pandas, Evaluate, Fastembed, Pylate
Appendix B Hyperparameters
Hyperparameter Value
LLM temperature inst: 0, reas: 11
freq_penalty 0
pres_penalty 0
think tokens QwQ: $20k$
HuatuoGPT: $8k$
RAG top_k $2\times$ num chunks
$k_{\text{RRF}}$ 6060
$k_{1}$ TF saturation
$b$ length norm.
avgdl avg. doc length
Table 6: Hyperparameter Configuration
Appendix C Datasets
C.1 Context Data Distributions
Token-length distributions per dataset are shown in Figures 6 and 7.
C.2 Context Data Preparation
We performed minimal cleaning and standardization to support retrieval and prompting, including merging notes by patient and stay, normalizing timestamps, adding note-type metadata, computing token counts for context segmentation, and filtering to human-verified subsets where applicable.
We describe the cleaning and normalization steps applied prior to indexing and prompting:
  • Cleaning and de-duplication of clinical notes; removal of template boilerplate where applicable.
  • Merging notes by patient and hospital stay; preserving note-type and section headers for downstream retrieval.
  • Datetime standardization and chronological ordering; normalization where timezones or partial timestamps occur.
  • Metadata augmentation (e.g., note type, encounter identifiers) to support Include DS vs Include All settings.
  • Tokenization and token-count computation at note- and patient-level for context segmentation.
Figure 6: MIMIC-III Context Distributions
Figure 7: MIMIC-IV Context Distributions
These details enable reproducibility of sampling and retrieval segmentation.
Dataset LLM-as-a-judge QA Pairs EHR-DS-QA Selene-8B 113 Prometheus-8x7B-v2.0 98 Qwen2.5-32B-Instruct 71 EHRNoteQA Selene-8B 180 Prometheus-8x7B-v2.0 220 Qwen2.5-32B-Instruct 42
Table 7: Disagreement instances where NLI Med Contradiction is high (>> 0.7) and LLM Correctness is also high (>> 0.7).
Dataset LLM-as-a-judge 1 LLM-as-a-judge 2 QA Pairs EHR-DS-QA Prometheus Qwen2.5-32B-Instruct 520 Prometheus Selene-8B 363 Qwen2.5-32B-Instruct Selene-8B 664 EHRNoteQA Prometheus Qwen2.5-32B-Instruct 690 Prometheus Selene-8B 236 Qwen2.5-32B-Instruct Selene-8B 519
Table 8: Judge-pair disagreements where the absolute difference in LLM Correctness $\geq$ 0.5.
Metric Splade BM25 LLM Correctness 67.30 69.34 LLM Completeness 65.41 67.14 LLM Faithfulness 47.64 46.23 NLI Med Entailment 52.68 54.35 NLI Med Contradiction 16.98 16.94 Bio BERTScore F1 79.75 79.67 METEOR 43.93 43.63
Table 9: Comparison of Different Sparse Retrievers
C.3 Prompt Formulations
We used a simple zero-shot prompt structure and tailored it explicitly for each task formulation:
  • Extractive: Request the model to answer by extracting the most relevant answer from the context.
  • Multiple-choice: Standard multiple choice prompt.
  • Open-ended: Focusing on open ended question answering favoring short, single sentence answers.
Below is the sample prompt for extractive tasks:
System message: Your input fields are: 1. medical_record (list[str]): List of patient notes (chronological order). 2. question (str): A question about the patients record. Your output fields are: 1. answer (str): Short single-sentence answer to the question. All interactions will be structured in the following way, with the appropriate values filled in. [[ ## medical_record ## ]] {medical_record} [[ ## question ## ]] {question} [[ ## answer ## ]] {answer} [[ ## completed ## ]] In adhering to this structure, your objective is: Given a patients medical record and a question, answer the question correctly and shortly.
C.4 Dataset Comparisons